# Analysis of Emotion in Rumour Threads on Social Media

### v1

@@ Abstract
Rumours in online social media pose significant risks to modern society, motivating the need for better understanding of how they develop. We focus specifically on the interface between emotion and rumours in threaded discourses, building on the surprisingly sparse literature on the topic which has largely focused on emotions within the original rumour posts themselves, and largely overlooked the comparative differences between rumours and non-rumours. In this work, we provide a comprehensive analytical emotion framework, contrasting rumour and non-rumour cases using existing NLP datasets to further understand the emotion dynamics within rumours. Our framework reveals several findings: rumours exhibit more negative sentiment and emotions, including anger, fear and pessimism, while non-rumours evoke more positive emotions; emotions are contagious in online interactions, with rumours facilitate negative emotions and non-rumours foster positive emotions; and based on causal analysis, surprise acts as a bridge between rumours and other emotions, pessimism is driven by sadness and fear, optimism by joy and love.
@@ Introduction
In part due to the ubiquity of edge devices like mobile phones, the major of the world’s population now has access to the internet. [^1] https://datareportal.com/reports/digital-2024-october-global-statshot This increasing ease of access and interaction through online social media has brought both opportunities and challenges. One significant challenge is the rapid spread of rumours. Rumours on online social media have become a major threat to society (Tian, Zhang, and Lau 2022; Zubiaga et al. 2015; Kochkina, Liakata, and Zubiaga 2018; Ma, Gao, and Wong 2017). The circulation of unsubstantiated rumours has impacted a large group of people, with consequences ranging from seeding skepticism and discrediting science, to endangering public health and safety. For example, during COVID-19, an Arizona man died, and his wife was hospitalized after ingesting a form of chloroquine in an attempt to prevent the disease. Additionally, 77 cell phone towers were set on fire due to conspiracy theories linking 5G networks to the spread of COVID-19 (Cui and Lee 2020). Recent advancements in Large Language Models (LLM) and generative AI have exacerbated this phenomenon, creating an urgent need to understand and better deal with rumours on social media (Chen and Shu 2024).
@@
Emotions have a strong influence on human behavior in both offline and online settings (Zollo et al. 2015; Herrando and Constantinides 2021; Ekman 1992). They shape the type of information users seek, how they process and remember it, and the judgments and decisions they make. Misinformation is often associated with high-arousal emotions such as anger, sadness, anxiety, surprise, and fear (Liu et al. 2024b). Rumours conveying these emotions are more likely to generate higher numbers of shares and exhibit long-lived, viral patterns (Pröllochs, Bär, and Feuerriegel 2021b). However, existing research on emotions in misinformation analysis is fragmented, focusing primarily on the emotions within the original rumour posts themselves and often overlooking the comparative differences between rumours and non-rumours (Ferrara and Yang 2015; Zollo et al. 2015). In this work, we address this gap by introducing a systematic analysis of various aspects of emotion, contrasting the emotional patterns in rumour and non-rumour content on social media. Using popular NLP rumour detection datasets, we provide new insights into the emotional dynamics of rumour and non-rumour data.
Data
We use three rumour datasets in this work, namely: PHEME (Zubiaga et al. 2015; Kochkina, Liakata, and Zubiaga 2018), Twitter15, and Twitter16 (Ma, Gao, and Wong 2017):
PHEME
Zubiaga et al. (2015) contains 6,425 tweet posts of rumours and non-rumours related to 9 events. To avoid using specific a priori keywords to search for tweet posts, PHEME used the Twitter (now X) steaming API to identify newsworthy events from breaking news and then selected from candidate rumours that met rumour criteria, finally they collected associated conversations and annotate them. They engaged journalists to annotate the threads. The data were collected between 2014 and 2015. The 9 events are split into two groups, the first being breaking news that contains rumours, including Ferguson unrest, Ottawa shooting, Sydney siege, Charlie Hebdo shooting, and Germanwings plane crash. The rest are specific rumours, namely Prince to play in Toronto, Gurlitt collection, Putin missing, and Michael Essien contracting Ebola.
Twitter 15
Liu et al. (2015) was constructed by collecting rumour and non-rumour posts from the tracking websites snopes.com and emergent.info. They then used the Twitter API to gather corresponding posts, resulting in 94 true and 446 false posts. This dataset further includes 1,490 root posts and their follow posts, comprising 1,116 rumours and 374 non-rumours.
Twitter 16
Similarly to Twitter 15, Ma et al. (2016) collected rumours and non-rumours from snopes.com, resulting in 778 reported events, 64% of which are rumours. For each event, keywords were extracted from the final part of the Snopes URL and refined manually—adding, deleting, or replacing words iteratively—until the composed queries yielded precise Twitter search results. The final dataset includes 1,490 root tweet posts and their follow posts, comprising 613 rumours and 205 non-rumours.
@@
Framework for Analyzing Emotion
In this section, we present our framework for analyzing emotion. We first establish a basic understanding of emotion polarity by determining the sentiment valence of each root tweet and comment. We then use multi-label emotion detection to predict the emotion categories associated with each post. Based on this data, we explore the interactive nature of emotions, by identifying common patterns in emotion transition pairs between temporally-adjacent posts. Finally we investigate the emotional trajectory within threads to understand how emotional intensity and type shift over time, by aggregating the predicted labels for posts at each time stamp in a given thread. As part of this, we contrast rumour with non-rumour threads, to gain a holistic understanding of emotional expression in rumours and non-rumours on Twitter.
Affective Computing: Automatic Emotion Detection
@@
Manually annotating emotions is both costly and time-consuming, so we use an LLM-based emotion detection model, EmoLLM (Liu et al. 2024a), which is specifically designed for sentiment analysis and emotion detection. The model was instruction-tuned on SemEval 2018 Task1 using a comprehensive emotion labeling scheme grounded in established theoretical frameworks. We prompt the model to perform Valence Ordinal Classification (V-oc), Emotion Classification (E-c), and Emotion Intensity regression (E-i). Detailed prompts are shown in Table 1.
Categorical Emotion Labeling Scheme
Numerous emotion label sets have been proposed (Ekman 1992; Plutchik 1980; Russell 1980). According to Ekman (1992); Plutchik (1980), certain emotions, such as joy, fear, and sadness, are considered more fundamental than others, both physiologically and cognitively. The Valence-Arousal-Dominance (VAD) model (Russell 1980) categorizes emotions within a three-dimensional space of valence (positivity-negativity), arousal (active-passive), and dominance (dominant-submissive). Inspired by Mohammad et al. (2018), we incorporate elements from both basic emotion theories and the VAD model, and further ground EmoLLM emotion predictions to develop the following emotion label schemes: (1) neutral or no emotion; (2) negative emotions: anger (also includes annoyance and rage), disgust (also includes disinterest, dislike, and loathing), fear (also includes apprehension, anxiety, and terror), pessimism (also includes cynicism, and no confidence), sadness (also includes pensiveness and grief); 3) positive emotions: joy (also includes serenity and ecstasy), love (also includes affection), optimism (also includes hopefulness and confidence), anticipation (also includes interest and vigilance), surprise (also includes distraction and amazement) and trust (also includes acceptance, liking, and admiration).
Emotion Polarity: Sentiment Valence
@@
To understand the basic emotion polarity expressed in rumour and non-rumour content, we begin with sentiment valence analysis. Sentiment valence aims to capture the overall emotional tone conveyed by a post, in terms of how positive or negative it is (Liu et al. 2024b). We frame the sentiment valence task as ordinal regression (Mohammad et al. 2018). As shown in Table 1, for a given tweet post, we classify it into one of seven ordinal levels of sentiment intensity, spanning varying degrees of positive and negative valence, that best represents the tweeter’s mental state. The tweet posts within a thread can be divided into two categories: root tweets, which are posted by the publisher, and follow posts, which include all subsequent replies under the root post. We begin by conducting sentiment valence analysis on each post within the thread conversation. For each category, we compute the mean sentiment valence to enable further investigation into the specific emotions associated with different sentiment valences over a thread.
Emotion Distribution
Following sentiment valence analysis, we then examine specific emotions and their distribution in rumour and non-rumour tweet posts. Motivated by the fact that a certain tweet might exhibit more than one emotion, we frame the task as multi-label emotion detection problem. As shown as V-oc in Table 1, given a tweet, we classify it into one of seven ordinal classes, corresponding to various levels of positive and negative sentiment intensity. To reduce noise from automatic emotion detectors, we take the top-three predicted emotions for each tweet. We then aggregate and plot the emotion distribution to provide an overview of dominant emotional trends across the rumour and non-rumour posts. Given that the follow posts make up the majority of the data compared to the root posts, we will focus on using follow posts in our next analysis.
Emotion Transitions
Emotions are contagious and highly interactive (Ferrara and Yang 2015). When publishers write tweets that convey their emotions, readers are likely to respond with emotional reactions of their own (Ferrara and Yang 2015; Zollo et al. 2015). In this part, we model this interactive nature of emotions in the form of emotion transition pairs, which are built from two chronologically-adjacent tweets. In each pair, the first element represents the emotion inferred from the initial content published at a given time, and the second element represents the emotion inferred from the reply content published immediately after. For example, if the first tweet exhibits joy trust and anticipation, and the second tweet shows anger, disgust and surprise, we form the pairs (joy, anger), (joy, disgust), (joy, surprise), (trust, surprise), (trust, surprise), (trust, disgust), (anticipation, anger), (anticipation, surprise) and (anticipation, disgust). We create transitions for all combinations of emotion pairs and explore the likelihood of emotion transition pairs occurring in rumour and non-rumour content. Exploring emotion transitions allows us to understand the emotional flow in social media conversations and uncover typical patterns of rumour and non-rumour content, and any differences between the two.
Emotion Trajectories
@@
PHEME root --0.25 --0.17 0.00 2602/2602 --0.21 --0.11 --0.39 7.75e-11 4.41e-11 629/629/629
PHEME follow --0.33 --0.26 6.47e-09 --0.35 --0.20 --0.39 0.03 8.38e-15
@@
Twitter15 follow --0.27 --0.06 1.65e-09 --0.24 --0.25 --0.30 0.16 0.21
Twitter16 root --0.18 --0.07 0.00 205/205 --0.11 --0.22 --0.30 1.35e-06 0.18 63/63/63
Twitter16 follow --0.31 --0.12 9.19e-06 --0.30 --0.36 --0.27 0.67 0.90
Table 2: Valence Ordinal Regression results for all datasets. root = root posts, follow = follow posts, Ru = rumour, Non = Non-rumour, T = True rumour, F = False rumour, U = Unverified rumour; $p$ values indicates significance of a one-tailed t-test.
Algorithm 1 PC Algorithm
1:Input: Data $\mathbf{X}$, significance level $\alpha$
2:Output: Completed Partially Directed Acyclic Graph (CPDAG)
4:Step 1: Skeleton Identification
5:for each pair of variables $(X,Y)$ in $G$ do
7:     if such a subset $S$ exists then
8:         Remove the edge $X-Y$ from $G$.
9:     end if
10:end for
11:Step 2: Edge Orientation
13:     if $Z\notin S$ for all separating sets $S$ for $X$ and $Y$ then
14:         Orient as $X\to Z\leftarrow Y$ (identify a collider).
15:     end if
16:end for
17:while possible do
18:     for each edge $(X-Y)$ in $G$ do
19:         if there exists a directed path $X\to\dots\to Z$ such that $Z-Y$ then
20:              Orient as $X\to Y$ (acyclicity rule).
21:         else if orienting $X-Y$ as $X\to Y$ creates a new v-structure then
22:              Orient as $X\to Y$ (v-structure rule).
23:         end if
24:     end for
25:end while
Causal Relationship of Emotions in Rumour & Non-Rumour Threads
To gain a deeper insight into the relationship between rumours and the emotions underlying them, we extend our analysis beyond statistical correlation by conducting a causal analysis. Specifically, we apply the Peter-Clark (PC) algorithm (Spirtes, Glymour, and Scheines 2000), a classical constraint-based causal discovery algorithm on the three merged datasets.
Uncovering causal relations between variables of interest is never an easy problem. Under the fundamental assumption of causal Markov condition that a variable is conditionally independent of all its non-effects given its direct cause, faithfulness ensures that the casual graph exactly encodes the independence and conditional independence relations among variables. These two assumptions allow us to infer causal relationships from observed statistical independencies, forming the cornerstone of constraint-based causal discovery methods.
The PC algorithm identifies causal relationships among the variables of interest, represented as a directed acyclic graph (DAG), by numerating the independence and conditional independence relationships. The algorithm consists of two main steps:
  1. 1.
  2. 2.
    Edge Orientation: After constructing the skeleton, edges are oriented by a set of predefined rules (Meek’s Rule (Meek 1997)) to avoid cycles and orient collider structures.
The complete PC algorithm is provided in algorithm 1. It returns a completed partially directed acyclic graph (CPDAG), which represents an equivalence class of causal graphs that are consistent with the observed data’s independence and conditional independence relations. In our implementation, we adopt the Fisher-z test (Fisher 1921) to infer the conditional independence relations.
Results and Discussion
In this section, we apply our framework to the collected data, and present experiment results and discuss our findings.
@@
We present the sentiment valence ordinal regression results in Table 2. The numbers are balanced by random down-sampling, i.e. rumour and non-rumour, true rumour and false rumour both have equal numbers of posts. As shown in the table, sentiment in rumour root posts and comments is significantly more negative than that in non-rumours across all datasets and settings ($p<0.05$). This means both publishers and commenters engaged in the thread exhibit a more negative mindset towards rumour content. Compared with rumour posts at the root level, comment posts exhibit more negative sentiment for all datasets. Additionally, we break down the rumour data into true, false, and unverified rumours according to their original labels in the dataset. Interestingly, we found that unverified content exhibits more negative sentiment compared to both true and false rumours in the PHEME dataset, as well as in the root posts of Twitter15 and the U vs. T setting in Twitter16. Given that sentiment is more negative in comments and they form the main part of the conversation, we conduct the following experiments using only comments.
@@
In order to further understand emotions expressed in rumour and non-rumour content, we present the emotion distribution results in Figures 1, 2 and 3. Overall, we observe a sharper distribution in emotions like anger, disgust, neutral, optimism, and joy. In the PHEME, Twitter15, and Twitter16 datasets, rumour posts tend to show more negative emotions such as anger, disgust, fear, and sadness, while non-rumour posts display more positive emotions like trust, optimism, joy and love. We present emotions statistics in Table 3.
@@
Table 5: Cumulative emotion regression coefficient across different datasets for rumour and non-rumour comments. Ang = Anger, Disg = Disgutst, Sad = Sadness, Pess = Pessimism, Neu = Neutral, Surp = Surprise, Antic = Anticipation, Opti = Optimism.

### v2

# An Analytical Emotion Framework of Rumour Threads on Social Media

@@ Abstract
Rumours in online social media pose significant risks to modern society, motivating the need for better understanding of how they develop. We focus specifically on the interface between emotion and rumours in threaded discourses, building on the surprisingly sparse literature on the topic which has largely focused on single aspect of emotions within the original rumour posts themselves, and largely overlooked the comparative differences between rumours and non-rumours. In this work, we take one step further to provide a comprehensive analytical emotion framework with multi-aspect emotion detection, contrasting rumour and non-rumour threads and provide both correlation and causal analysis of emotions. We applied our framework on existing widely-used rumour datasets to further understand the emotion dynamics in online social media threads. Our framework reveals that rumours trigger more negative emotions (e.g., anger, fear, pessimism), while non-rumours evoke more positive ones. Emotions are contagious—rumours spread negativity, non-rumours spread positivity. Causal analysis shows surprise bridges rumours and other emotions; pessimism comes from sadness and fear, while optimism arises from joy and love.
@@ Introduction
In part due to the ubiquity of edge devices like mobile phones, the major of the world’s population now has access to the internet. [^1] https://datareportal.com/reports/digital-2024-october-global-statshot This increasing ease of access and interaction through online social media has brought both opportunities and challenges. One significant challenge is the rapid spread of rumours. Rumours on online social media have become a major threat to society (Tian, Zhang, and Lau 2022; Zubiaga et al. 2015a; Kochkina, Liakata, and Zubiaga 2018; Ma, Gao, and Wong 2017). The circulation of unsubstantiated rumours has impacted a large group of people, with consequences ranging from seeding skepticism and discrediting science, to endangering public health and safety. For example, during COVID-19, an Arizona man died, and his wife was hospitalized after ingesting a form of chloroquine in an attempt to prevent the disease. Additionally, 77 cell phone towers were set on fire due to conspiracy theories linking 5G networks to the spread of COVID-19 (Cui and Lee 2020). Recent advancements in Large Language Models (LLM) and generative AI (OpenAI 2023; Anthropic 2024; Dubey et al. 2024) have exacerbated this phenomenon, creating an urgent need to understand and better deal with rumours on social media (Chen and Shu 2024).
@@
Emotions have a strong influence on human behavior in both offline and online settings (Zollo et al. 2015; Herrando and Constantinides 2021; Ekman 1992). They shape the type of information users seek, how they process and remember it, and the judgments and decisions they make. Misinformation is often associated with high-arousal emotions such as anger, sadness, anxiety, surprise, and fear (Liu et al. 2024b). Rumours conveying these emotions are more likely to generate higher numbers of shares and exhibit long-lived, viral patterns (Pröllochs, Bär, and Feuerriegel 2021b).
Existing research on emotions in rumour analysis can broadly be categorized into two strands: (1) studies that leverage emotional signals to assist rumour detection (Ferrara and Yang 2015; Zollo et al. 2015; Liu et al. 2024b), and (2) studies that analyze emotional dynamics within rumour threads to understand rumour propagation pattern (Pröllochs, Bär, and Feuerriegel 2021b, a; Wang et al. 2021; Marino, Benitez-Baleato, and Ribeiro 2024) and emotion contagion pattern (Kramer, Guillory, and Hancock 2014; Coviello et al. 2014).
However, much of this work remains fragmented and has several systematic limitations. They often focus narrowly on the rumour posts with limited emotional aspects considered, and they rarely compare rumour and non-rumour threads. Moreover, most existing studies explore correlations between emotion and rumours—insights of causal relationship into how emotions affect rumours are lacking. In this work, we address this gap by taking one step further to provide a comprehensive emotion analysis framework for rumour threads in online social media. Our framework provides a wide range of analysis of emotions from basic emotion polarity, emotion distribution to emotion patterns like transitions, trajectories and causal relationship of emotions within online rumour threads.
Our contribution can be summarized as follows:
  • We go beyond prior work that focuses on a single or few emotion dimension by performing automatic, multi-aspect emotion detection and analysis, offering broader coverage of emotional signals in rumour threads.
  • We contrast emotional patterns between rumour and non-rumour instances and provide both correlation and causal insights in the hope to support future rumour detection systems.
  • We conduct analysis with our framework in three widely-used rumour datasets from online social media, demonstrating the feasibility of the framework at scale.
@@
Related Work
The definition of rumour is generally complicated and varies from one publication to another. Some early work treated rumour as information that is false (Cai, Wu, and Lv 2014). Recent definitions of rumours are “unverified and instrumentally relevant information statements in circulation” (DiFonzo and Bordia 2007) and “unverified information at the time of the posting”. This definition also aligns with the concept in recent work (Zubiaga et al. 2018, 2015a; Tian, Zhang, and Lau 2022) and the Oxford English Dictionary, which defines the rumour as “an unverified or unconfirmed statement or report circulating in a community”. [^2] https://www.oed.com/dictionary/rumour_n?tab=meaning_and_use
Existing research highlights the significant role of emotions in understanding general misinformation, mostly fake news. Research has found relationships exist between negative sentiment and fake news, and between positive sentiment and genuine news (Zaeem et al. 2020). Fake news also expresses a higher level of overall emotion, negative emotion, and anger than real news (Zhou, Tao, and Zhang 2022). Negative emotions like sadness and anger can serve as indicators of misinformation (Prabhala and Bose 2019). The role of emotions in rumours has been recognized since the Second World War, reflecting the interactive and community-driven nature of rumour spreading. Knapp’s taxonomy (Knapp 1944) of rumours categorizes them into three types, each deeply embedded with emotions: (1) ‘pipedream’ rumours, which evoke wishful thinking; (2) ‘bogy’ rumours, which heighten anxiety or fear; and (3) ‘wedge-driving’ rumours, which incite hatred. This taxonomy underscores how rumours are inherently embedded with emotional undercurrents.
Recent research on emotion in rumours largely focuses on their role in spreading behaviour. Some studies have used questionnaires to gather participants’ reactions to specific rumours (Zhang et al. 2022; Rijo and Waldzus 2023; Ali et al. 2022), while others have employed cascade size and lifespan as indicators (Pröllochs, Bär, and Feuerriegel 2021b, a). Key findings of such work include: rumours conveying anticipation, anger, trust, or offensiveness tend to generate more shares, have longer lifespans, and exhibit higher virality (Pröllochs, Bär, and Feuerriegel 2021b). Additionally, false rumours containing a high proportion of terms reflecting positive sentiment, trust, anticipation, anger, or condemnation are more likely to go viral (Solovev and Pröllochs 2022; Pröllochs, Bär, and Feuerriegel 2021a). However, existing research has notable gaps: it often focuses on isolated and limited aspects of emotions in rumours, primarily identifies correlations rather than causality, and tends to examine rumour data alone. To address these gaps, we aim to propose a comprehensive emotional analytical framework that integrates multiple emotion-related tasks, contrasting rumour and non-rumour content, providing a more comprehensive way to analyze rumour and non-rumour threads with the aim to enhance our understanding of emotions and, ultimately, to provide insights to facilitate rumour detection and analysis in online social media.
Data
In this section, we provide details of the rumour datasets used for analysis of emotion in rumour threads on social media. We adopt 3 widely used rumour datasets: PHEME (Zubiaga et al. 2015a; Kochkina, Liakata, and Zubiaga 2018), Twitter15, and Twitter16 (Ma, Gao, and Wong 2017). We introduce their details as follows:
PHEME
Zubiaga et al. (2015a) contains 6,425 tweet posts of rumours and non-rumours related to 9 events. To avoid the bias introduced by using a priori keywords—i.e., identifying rumours based on prior knowledge of specific events or predefined keywords rather than discovering them dynamically, PHEME used the Twitter (now X) streaming API to identify newsworthy events from breaking news. First, they collected candidate rumourous stories signaled by highly retweeted tweets linked to newsworthy current events. Next, journalists on the research team manually examined a subset of samples and selected those that met established rumour criteria (Zubiaga et al. 2015b) and identified the specific tweets that introduced them. Finally, they collected conversations (threads) associated with these rumour-introducing tweets for further analysis.
The data were collected between 2014 and 2015 and cover 9 events, divided into two groups: breaking news events likely to spark multiple rumours, and specific rumours identified a priori. The first group includes five cases—Ferguson unrest, Ottawa shooting, Sydney siege, Charlie Hebdo shooting, and the Germanwings plane crash. The second group comprises four specific rumours: Prince to play in Toronto, Gurlitt collection, Putin missing, and Michael Essien contracting Ebola.
Twitter 15
Liu et al. (2015) built the dataset by crawling two rumour verification websites (Snopes.com and Emergent.info), resulting in 2,299 candidate stories posted up to March 2015. To gather relevant tweets for each story, the authors formulated keyword-based queries combining subjects/objects with potential actions, and submitted them directly to Twitter’s search interface to retrieve historical tweets. Researchers manually verified the results through sampling. To balance the dataset with additional true stories, the authors also used Twitter’s 1% streaming API to identify newsworthy and credible events. This process produced 421 true and 421 false events. To gather conversation threads, Twitter15 includes 1,490 root posts and their associated comment posts, comprising 1,116 rumours and 374 non-rumours in the final dataset.
Twitter 16
Similarly to Twitter 15, Ma et al. (2016) collected rumours and non-rumours from Snopes.com. The authors identified 778 reported events between March and December 2015, of which 64% were rumours. For each event, they extracted keywords from the final part of the Snopes URL and refined manually to ensure that the resulting queries to Twitter search interface return precise results. The final dataset includes 1,490 root tweet posts and their associated comment posts, comprising 613 rumours and 205 non-rumours conversations.
Data Structure and Labels
All tweet posts within a thread can be divided into two categories: root tweets, which are posted by the publisher, and comment posts, which include all subsequent replies under the root post. All datasets provide a binary label—rumour or non-rumour at the conversation thread level. In addition, rumours are annotated with one of three extra labels: True, False, or Unverified, indicating the final truth status of the rumour (Zubiaga et al. 2015a; Liu et al. 2015; Ma et al. 2016). All rumours start off in an Unverified state. A rumour is labeled True if it is ultimately confirmed to be genuine, and False if it is misinformation. If the truth status remains unclear at the time of dataset creation, the rumour thread remains Unverified.
Automatic Emotion Information Annotation
@@
Manually annotating emotions is both costly and time-consuming. With the advancement of natural language processing (NLP), researchers adopt emotion detection models to label affective information automatically at scale. In this work, we use an emotional large language model and annotation tools, EmoLLM (Liu et al. 2024a), to conduct automatic emotion annotation. [^3] EmoLLMs contains a series of emotional large language models based on LLaMA (Touvron et al. 2023), we used EmoLLaMA-chat-13B in our experiment. EmoLLM was instruction-tuned on SemEval 2018 Task1: Affect in Tweets using a comprehensive emotion labeling scheme grounded in established theoretical frameworks (Mohammad et al. 2018). We annotate data with EmoLLM across three tasks: Valence Ordinal Classification (V-oc), Emotion Classification (E-c), and Emotion Intensity regression (E-i). Detailed prompts are shown in Table 1.
Emotion Polarity: Sentiment Valence (V-oc)
@@
To understand the basic emotion polarity expressed in rumour and non-rumour content, we begin with sentiment valence analysis based on V-oc. Sentiment valence aims to capture the overall emotional tone conveyed by a post, in terms of how positive or negative it is (Liu et al. 2024b). As shown in Table 1, for a given tweet post, we classify it into one of 7 ordinal levels of sentiment intensity, spanning varying degrees of positive and negative valence, that best represents the tweeter’s mental state.
Categorical Emotion Classification Scheme (E-c)
Numerous emotion labeling schemes have been proposed (Ekman 1992; Plutchik 1980; Russell 1980). According to Ekman (1992); Plutchik (1980), certain emotions, such as joy, fear, and sadness, are considered more fundamental than others, both physiologically and cognitively. The Valence-Arousal-Dominance (VAD) model (Russell 1980) categorizes emotions within a three-dimensional space of valence (positivity-negativity), arousal (active-passive), and dominance (dominant-submissive). Inspired by Mohammad et al. (2018), we incorporate elements from both basic emotion theories and the VAD model, and further ground EmoLLM emotion classifications to develop the following emotion label schemes: (1) neutral or no emotion; (2) negative emotions: anger (also includes annoyance and rage), disgust (also includes disinterest, dislike, and loathing), fear (also includes apprehension, anxiety, and terror), pessimism (also includes cynicism, and no confidence), sadness (also includes pensiveness and grief); 3) positive emotions: joy (also includes serenity and ecstasy), love (also includes affection), optimism (also includes hopefulness and confidence), anticipation (also includes interest and vigilance), surprise (also includes distraction and amazement) and trust (also includes acceptance, liking, and admiration).
Emotion Intensity Regression (E-i)
Capturing the full spectrum of emotions in online texts requires moving beyond simple emotion classification to understanding its intensity. Our expressions inherently convey varying degrees of feeling such as being very angry, slightly sad, absolutely joyful, etc. Quantifying this emotional intensity offers valuable insights with applications spanning commerce, public health initiatives, intelligence analysis, and social welfare (Mohammad and Bravo-Marquez 2017). The task can dates back to early work in the WASSA-2017 Shared Task on Emotion Intensity (Mohammad and Bravo-Marquez 2017), this task has remained a key challenge in affective computing and continuing in the SemEval-2018 Task 1: Affect in Tweets (Mohammad et al. 2018) and recent SemEval-2025 Task 11: Bridging the Gap in Text-Based Emotion Detection (Muhammad et al. 2025). A great advantage of the datasets associated with these tasks is their reliance on the Best-Worst Scaling annotation (Mohammad and Bravo-Marquez 2017), which lead to more reliable fine-grained intensity scores.
Human Evaluation
Although the performance of EmoLLM on emotion tasks was validated in  Liu et al. (2024a), we further evaluate its performance in our datasets with human evaluation. Specifically, we randomly sampled 50 instances from PHEME, Twitter15, and Twitter16, and conducted manual annotation for V-oc, E-c, and E-i tasks with 3 annotators specialized in NLP. Annotators were provided with clear annotation guidelines and received training before annotation. On average, we achieved a final annotator agreement of Cohen’s kappa of 0.50 and Pearson’s correlation of 0.51. The full annotation guideline is included in the Appendix.
@@
PHEME root --0.25 --0.17 0.01 2602/2602 --0.21 --0.11 --0.39 7.75e-11 4.41e-11 629/629/629
PHEME comment --0.33 --0.26 6.47e-09 --0.35 --0.20 --0.39 0.03 8.38e-15
@@
Twitter15 comment --0.27 --0.06 1.65e-09 --0.24 --0.25 --0.30 0.16 0.21
Twitter16 root --0.18 --0.07 0.01 205/205 --0.11 --0.22 --0.30 1.35e-06 0.18 63/63/63
Twitter16 comment --0.31 --0.12 9.19e-06 --0.30 --0.36 --0.27 0.67 0.90
Table 2: Valence Ordinal Classification results for all datasets. root = root posts, comment = comment posts to the root posts, Ru = rumour, Non = Non-rumour, T = True rumour, F = False rumour, U = Unverified rumour; $p$ values indicates significance of the one-tailed t-test.
Framework for Analyzing Emotions
In this section, we present our framework for analyzing emotion. We first establish a basic understanding of emotion polarity by determining the sentiment valence of each root and comment tweet. Then we apply multi-label emotion detection to predict the emotion categories associated with each post. Based on this data, we explore the interactive nature of emotions, by identifying common patterns in emotion transition pairs between temporally-adjacent posts. Finally we investigate the emotional trajectory within threads to understand how emotional intensity and type shift over time, by aggregating the predicted labels for posts at each time stamp in a given thread. As part of this, we contrast rumour with non-rumour threads, to gain a holistic understanding of emotional expression in rumours and non-rumours on Twitter.
@@
We begin by conducting sentiment valence analysis on each post within the thread. For each category, we compute the mean sentiment valence to enable further investigation into the specific emotions associated with different sentiment valences over a thread. We present the sentiment valence ordinal regression results in Table 2. The numbers are balanced by random down-sampling, i.e. rumour and non-rumour, true rumour and false rumour both have equal numbers of posts. As shown in the table, sentiment in rumour root posts and comments is significantly more negative than that in non-rumours across all datasets and settings ($p<0.05$). This means both publishers and commenters engaged in the thread exhibit a more negative mindset towards rumour content. Compared with rumour posts at the root level, comment posts exhibit more negative sentiment for all datasets. Additionally, we break down the rumour data into true, false, and unverified rumours according to their original labels in the dataset. Interestingly, we found that unverified content exhibits more negative sentiment compared to both true and false rumours in the PHEME dataset, as well as in the root posts of Twitter15 and the U vs. T setting in Twitter16. Given that sentiment is more negative in comments and they form the main part of the conversation, we conduct the following experiments using only comment posts.
@@
Following sentiment valence analysis, we then examine specific emotions and their distribution in rumour and non-rumour tweet comment posts. Motivated by the fact that a certain tweet might exhibit more than one emotion, we frame the task as multi-label emotion detection problem. As shown as E-c in Table 1, given a tweet post, we classify it into one or more emotions in 11 classes. We take the top3 predicted emotions as dominant ones for each post. We then aggregate and plot the emotion distribution to provide an overview of dominant emotional trends across the rumour and non-rumour posts. Given that the comment posts make up the majority of the data compared to the root posts, we focus on using comment posts in our following analysis. We present the emotion distribution in the comments in Figures 1, 2 and 3. Overall, we observe a sharper distribution in emotions like anger, disgust, neutral, optimism, and joy. Generally, in PHEME, Twitter15, and Twitter16 datasets, comments in rummour threads tend to show more negative emotions such as anger, disgust, fear, and sadness, while comments in non-rumour threads display more positive emotions like trust, optimism, joy and love. We present emotion statistics over the comments in rumour and non-rumour threads in Table 3.444It is important to note that while this represents a general trend observed across the datasets, there are exceptions. For instance, Figure 1 reveals that non-rumour threads in PHEME exhibit higher instances of anger and disgust compared to rumour threads.
@@
Table 5: Cumulative emotion regression coefficient across different datasets for rumour and non-rumour comments. Ang = Anger, Disg = Disgust, Sad = Sadness, Pess = Pessimism, Neu = Neutral, Surp = Surprise, Antic = Anticipation, Opti = Optimism. Larger value indicates a more rapid growth rate.
@@
Emotions are contagious and highly interactive (Ferrara and Yang 2015). When publishers write tweets that convey their emotions, readers are likely to respond with emotional reactions of their own (Ferrara and Yang 2015; Zollo et al. 2015). In this part, we model this interactive nature of emotions in the form of emotion transition pairs, which are built from two chronologically-adjacent tweet posts. In each pair, the first element represents the emotion inferred from a tweet posted at a given time, and the second element represents the emotion inferred from the tweet posted immediately after the former tweet. For example, if the first tweet exhibits joy trust and anticipation, and the second tweet shows anger, disgust and surprise, we form the pairs (joy, anger), (joy, disgust), (joy, surprise), (trust, surprise), (trust, surprise), (trust, disgust), (anticipation, anger), (anticipation, surprise) and (anticipation, disgust). We create transitions for all combinations of emotion pairs and explore the likelihood of emotion transition pairs occurring in rumour and non-rumour content. Exploring emotion transitions allows us to understand the emotional flow in social media conversations and uncover typical patterns of rumour and non-rumour content, and any differences between the two.